A transverse wave is applied to a string whose mass per unit length is 3×10 raise to power -2kgmper -1.if the string is under a tension of 120N, calculate the speed of propagation of wave?
1
Expert's answer
2020-10-16T10:59:47-0400

Given

Linear density is "\mu=3.0\times10^{-2} kg/m"

Tension is "T=120 N"

The expression for speed of propagation of wave is

"v=\sqrt \frac{T}{\mu}= \sqrt \frac{120 N}{3.0\times10^{-2} kg/m}=63 m/s"